OpenStudy (anonymous):

Use the grouping method to factor the polynomial below completely. 3x3 + 9x2 + 4x + 12 A. (3x2 + 3)(x + 3) B. (3x2 + 3)(x + 4) C. (3x2 + 4)(x + 4) D. (3x2 + 4)(x + 3)

Parth (parthkohli):

Take the common factor of the first two terms, and then the second two terms. Put them outside the bracket. Both the brackets should have the same values. Do you get it?

OpenStudy (anonymous):

noooooooooooo realkly dont

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\rightarrow 3x^3 + 9x^2 = 3x^2(x + 3)}\) \(\Large \color{MidnightBlue}{\rightarrow 4x + 12 = 4(x + 3) }\)

Parth (parthkohli):

Can you do it now?

OpenStudy (anonymous):

3x3 + 9x2 + 4x + 12 (3x^3+9x^2)+(4x+12) 3x^2(x+3)+4(x+3) (3x^2+4)(x+3)

Parth (parthkohli):

Use the distibutive property to combine.

OpenStudy (anonymous):

B. (3x2 + 3)(x + 4)

Parth (parthkohli):

The fourth one.